OpenStudy (anonymous):

kevin ran 3 miles less than steve ran .the sum of their distance is 21 miles.how far did steve run? the domain of the solution (0,10.5,12,18) uses every step

OpenStudy (anonymous):

K= S-3 K+S=21

OpenStudy (anonymous):

where did u get k

OpenStudy (anonymous):

k stands for kevin

OpenStudy (anonymous):

he ran 3 miles less than steve

OpenStudy (anonymous):

ok but i have to use all the steps 1.define the varbials 2.write algebraic expression for the distance each person ran 3.write an equation that can help u solve the problem 4.use the given replacement sets to solve the equation 5.check your answer

OpenStudy (anonymous):

1. Variable K and S (K= Kevin, S=Steve) 2. I did that 3.

OpenStudy (anonymous):

If you add both equations together you can solve them. K = S-3 K = -S +21 + --------------- 2K = 18 K = 9 9 = S - 3 S =12

OpenStudy (anonymous):

but it needs to be 21 not 12

OpenStudy (anonymous):

12+9 = 21